OpenStudy (anonymous):

Suppose that {f(x) = {-1}/{{-7}x^{6}}}. Evaluate each of the following: f '(3) f '(-1)

OpenStudy (anonymous):

\[\frac{ -1 }{ -7x^6 }\]

OpenStudy (anonymous):

i got -0.0192043896 for f'(3) and 42 for f'(-1) but it says it's wrong.

OpenStudy (loser66):

yours is too many { }. I don't get . is it \(f(x)= \dfrac{-1}{-7x^6}\) or \(f(x)=\dfrac{-1}{(-7x)^6}\)

OpenStudy (anonymous):

the first one

OpenStudy (loser66):

so, the question is about derivative,right? what is f'(x)?

OpenStudy (anonymous):

yes. i got 42x^-7, but idk if it's right

OpenStudy (loser66):

nope, -7 from denominator , it's still there yours = \(\dfrac{x^{-6}}{7}\)

OpenStudy (loser66):

f'(x) =\(\dfrac{-6x^{-7}}{7}\)

OpenStudy (anonymous):

got it.. thanks